OpenStudy (anonymous):

A designer is creating shirts that each have 12 buttons. She bought a container of 115 buttons. What are the possible numbers of shirts she can make?

OpenStudy (anonymous):

The designer can create \[\lfloor \frac{ 115 }{ 12 } \rfloor = 9 shirts\] at most.

OpenStudy (anonymous):

Need an algebraic equation for "numbers of shirts she can make"

OpenStudy (anonymous):

You can write in this way. N is the numbers of shirts she can make. \[0 \le N \le 9\]